Check whether the function f, given by
f(x) = cos x - sin x has an extreme value in the interval (-à ƒ ƒ  €/2,à ƒ ƒ  €/2).
1
Expert's answer
2021-10-14T18:15:08-0400

f(x) = cos x - sin x

Differentiating with respect to x

f'(x) = - sin x - cos x

f''(x) = - cos x + sin x

For extreme values of f(x) ,

f'(x) = 0

=> - sin x - cos x = 0

=> sin x = - cos x

=> tan x = -1

=> tan x = tan ("-\frac{π}{4})"

=> x = nπ "-\frac{π}{4}" Where n is an integer

Since x "\in(-\frac{π}{2}, \frac{π}{2})" the only value of n is 0.

So x = "-\frac{π}{4}"

f''("-\frac{π}{4}) = -cos(-\frac{π}{4})+sin(-\frac{π}{4})"

= "-cos(\frac{π}{4})-sin(\frac{π}{4})"

= "-\frac{1}{\sqrt{2}} -\frac{1}{\sqrt{2}}"

= "-\frac{2}{\sqrt{2}}"

= "-\sqrt{2}"

So f''("-\frac{π}{4})" is negative.

Therefore f(x) has a maximum value at x = "-\frac{π}{4}" and maximum value is "f(-\frac{π}{4}) = cos(-\frac{π}{4})-sin(-\frac{π}{4})"

= "\frac{1}{\sqrt{2}} +\frac{1}{\sqrt{2}}"

="\sqrt{2}"

Answer:

Yes , there is a extreme value of f(x) = cos x - sin x at x = "-\frac{π}{4}" and it is a maximum value equating "\sqrt{2}"